# A preorder on the set of links defined via orbifolds

### v1

@@ 8.2. Questions on bridge numbers
The answer to Question 8.4 is positive when $L$ is a Montesinos knot by Corollary 1.7. It is also true when $L^{\prime}$ has at most three bridges, and hence when $G^{\mathrm{orb}}(L^{\prime})$ is finite (by the orbifold theorem).
@@
Question 8.4 is also positive for torus knots.
@@
Let us assume that $K_{1}$ is a $(p_{1},q_{1})$-torus knot and $K_{2}$ a $(p_{2},q_{2})$-torus knot with $3\leq p_{1}<q_{1}$, $3\leq p_{2}<q_{2}$, $p_{1},q_{1}$ coprime, and $p_{2},q_{2}$ coprime. Furthermore, if $p_{i}=3$, then $q_{i}\geq 7$. Since $K_{1}\succeq K_{2}$ there is an epimorphism $\varphi\colon G^{\mathrm{orb}}(K_{1})\twoheadrightarrow G^{\mathrm{orb}}(K_{2})$ which, by the proof of Theorem 1.5(4), sends the infinite cyclic center $Z_{1}$ of $G^{\mathrm{orb}}(K_{1})$ into the infinite cyclic center $Z_{2}$ of $G^{\mathrm{orb}}(K_{2})$. The quotients $G(K_{1})/Z_{1}\cong\pi_{1}^{\mathrm{orb}}(S^{2}(2,p_{1},q_{1}))=T(2,p_{1},q_{1})$ and $G(K_{2})/Z_{2}\cong\pi_{1}^{\mathrm{orb}}(S^{2}(2,p_{2},q_{2}))=T(2,p_{2},q_{2})$ are hyperbolic triangle groups since $\frac{1}{2}+\frac{1}{p_{i}}+\frac{1}{q_{i}}<1$ for $i=1,2$. In particular, they are centerless, and thus the epimorphism $\varphi$ induces an epimorphism $\bar{\varphi}\colon T(2,p_{1},q_{1})\twoheadrightarrow T(2,p_{2},q_{2})$. Like in the proof of Proposition 6.6, it follows from the presentation of the triangle group $T(2,p_{1},q_{1})$ that each image $\bar{\varphi}(x)$, $\bar{\varphi}(y)$ and $\bar{\varphi}(z)$ is not trivial, otherwise the image $\bar{\varphi}(T(2,p_{1},q_{1}))$ would be trivial. Therefore, the non-trivial elements $a=\bar{\varphi}(x)$ and $b=\bar{\varphi}(y)$ verifies $a^{2}=1$ and $b^{p^{\prime}_{1}}=1$ with $p^{\prime}_{1}$ dividing $p_{1}$. The elements $a$ and $b$ generate the image $\bar{\phi}(T(2,p_{1},q_{1}))=T(2,p_{2},q_{2})$ which is a discrete subgroup of $\mathit{PSL}(2,\mathbb{R})$ since $T(2,p_{2},q_{2})$ is a hyperbolic triangle group. As in the proof of Proposition 6.6, up to taking suitable powers $u=a^{\pm 1}$ and $v=b^{k}$, $k$ coprime to $p^{\prime}_{1}$ to normalize the matrix representatives of $u$ and $v$ in $SL(2,\mathbb{R})$, it follows that at least one of Cases (I)–(VII) in [42, Theorem 2.3] holds. Since the triangle group $T(2,p_{2},q_{2})$ is co-compact, Case (II) is impossible by [42, Figures 2 and 3]. Case (III) cannot happens, otherwise $0=\operatorname{tr}(u)=\operatorname{tr}(v)$ and $\bar{\varphi}(T(2,p_{1},q_{1}))=T(2,\ell,n)$ with $\operatorname{tr}(u)=\operatorname{tr}(v)=\cos(\pi-\frac{\pi}{\ell})$ by [42, Section 3]. Then $\ell=2$, which contradicts $\frac{1}{2}+\frac{1}{\ell}+\frac{1}{n}<1$ because $T(2,\ell,n)=T(2,p_{2},q_{2})$ is a hyperbolic triangle group. If Case (IV) happens, then $u$ and $v$ generate a hyperbolic triangle group $T(2,3,n)$ by [42, Section 3]. Since $T(2,3,n)=T(2,p_{2},q_{2})$ with $3\leq p_{2}<q_{2}$ and $p_{2}$ coprime to $q_{2}$, it follows that $p_{2}=3$ and $q_{2}=n$. Therefore,
@@
For a prime link with infinite orbifold group, a positive answer to Question 8.4 would follow from a positive answer to the following question which is a stronger version of the well-known meridional rank conjecture by Cappell and Shaneson [1, Problem 1.11] about whether there is an equality between the bridge number $b(L)$ and the meridional rank $m(L)$ of a link $L$.
Question 8.7.
@@
For a link $L$ as in Question 8.7, $\Sigma_{2}(L)$ is aspherical by Lemma 3.1(3). Therefore, a torsion element in $G^{\mathrm{orb}}(L)$ has order $2$ and is the image of a meridian of $L$.
@@
It follows from results in [2] and [3] that the answer to Question 8.7 is positive for some special classes of links including Montesinos links, twisted links (see [2] for the definition), and arborescent links associated to bipartite trees with even weights. In these articles, it is shown that for such a link, a lower bound on the meridional rank obtained via Coxeter quotients of the link group matches the bridge number of the link. A Coxeter quotient of a link group $G(L)$ is a Coxeter group $C$ with an epimorphism from $G(L)$ onto $C$ which sends the meridians of $G(L)$ to reflections in $C$. The reflection rank of a Coxeter group $C$ is the minimal number of reflections needed to generate $C$. The Coxeter rank $\mathrm{cox}(L)$ of a link $L$ is the maximum of the reflection ranks of the Coxeter quotients of $G(L)$. For a prime link with infinite orbifold group, the inequalities
@@
The analog of Question 8.4 for the domination relation $\geqslant$ is also open for knots in $S^{3}$. It is not true for links because there are links with homeomorphic complements, but with different meridional rank and bridge number. For example, the untwisted Whitehead link $W_{1}$ has meridional rank $2$ and bridge number $2$ while the twisted Whitehead link $W_{n}$ obtained by $n$ full twists around one component of $W_{1}$ has meridional rank and bridge number $3$ if $|n|\geq 2$. For links, the analog of Question 8.4 makes sense for the $1$-domination relation $\geqslant_{1}$. In fact, a positive answer to Question 8.4 would show that $L\geqslant_{1}L^{\prime}$ implies $b(L)\geq b(L^{\prime})$.
8.3. Questions on the inequalities of the genus or the volume
Fifty years ago, Jonathan Simon [61] asked whether the domination relation $K_{1}\geq K_{2}$ between two knots implies the genus inequality $g(K_{1})\geq g(K_{2})$. The answer to this question also remains open for the $1$-domination relation. This property is definitely false for the $\pi$-domination. In [5, Proposition 4.1], for each positive odd integer $n$, we construct a hyperbolic knot $K_{n}$ of genus $2$ which admits a symmetric union presentation whose partial knot is the $(2,2n+1)$-torus knot $T(2,2n+1)$. Therefore, $K_{n}\succeq T(2,2n+1)$, but $T(2,2n+1)$ is of genus $n$. It is worth noting that, for $n>2$, $K_{n}\not\geqslant T(2,2n+1)$ since the Alexander polynomial of $K_{n}$ is not divisible by that of $T(2,2n+1)$ because $2n=\deg\Delta_{T(2,2n+1)}>4\geq\deg\Delta_{K_{n}}$.
The volume of a closed orientable $3$-dimensional orbifold $\mathcal{O}$ is defined as the sum of the volumes of the hyperbolic pieces in the geometric decomposition of $\mathcal{O}$. Then, for the closed $3$-dimensional orbifold $\mathcal{O}(L)$ with underlying space $S^{3}$ and singular locus a link $L$ whose branching index is $2$, one has the equality $\mathrm{vol}(\mathcal{O}(L))=\mathrm{vol}(\Sigma_{2}(L))/2$. Theorems 1.5 and 1.8 raise the following question.
Question 8.11.
Let $L_{1},L_{2}\subset S^{3}$ be links. Does $L_{1}\succeq L_{2}$ imply $\mathrm{vol}(\mathcal{O}(L_{1}))\geq\mathrm{vol}(\mathcal{O}(L_{2}))$?
Theorem 1.5 implies that the answer to Question 8.11 is positive for $L_{1}$ being a $2$-bridge link, a Montesinos link, or a Seifert link with non-vanishing determinant. A positive answer to Question 8.11 would imply that an arborescent link $\pi$-dominates only arborescent links, thus greatly improving Theorem 1.8. In general, Question 8.11 is widely open.
As above for a link $L\subset S^{3}$, the volume $\mathrm{vol}(E(L))$ is defined as the sum of the volumes of the hyperbolic pieces in the geometric decomposition of the exterior $E(L)$ of $L$. In Question 8.11, one cannot replace the volume $\mathrm{vol}(\mathcal{O}(L))$ of the orbifold $\mathcal{O}(L)$ by the volume $\mathrm{vol}(E(L))$. By Proposition 5.5, a link $L^{\prime}$ with $\det L^{\prime}=0$ $\pi$-dominates every $2$-bridge links $L(\frac{p}{q})$. The volumes of the exteriors of $2$-bridge links are unbounded, therefore there are infinitely many $2$-bridge links $L(\frac{p}{q})$ such that $\mathrm{vol}(E(L^{\prime}))<\mathrm{vol}(E(L(\frac{p}{q})))$.
With respect to the volume of the exterior of a link, it is more natural to consider the domination relation $\geqslant$.
Question 8.12.
Let $L_{1},L_{2}\subset S^{3}$ be links. Does $L_{1}\geqslant L_{2}$ imply $\mathrm{vol}(E(L_{1}))\geq\mathrm{vol}(E(L_{2}))$?
The answer to this question is open even for the $1$-domination relation $\geqslant_{1}$. It is true for $K_{1}$ being a $2$-bridge knot or a torus knot, but it is still open for a Montesinos knot. For a $2$-bridge knot $K_{1}$, any epimorphism $G(K_{1})\twoheadrightarrow G(K_{2})$ is induced by a map of non-zero degree, see [4, Corollary 1.3]. Then the inequality follows from the fact that $\mathrm{vol}(E(K_{1}))\geq k\mathrm{vol}(E(K_{2}))$, where $k\in\mathbb{Z}_{>0}$ is the degree of the map. If $K_{1}$ is a torus knot and $K_{1}\geqslant K_{2}$, then $K_{2}$ is a torus knot (see [4, Section 3.1] for example). Therefore, $\mathrm{vol}(E(K_{1}))=\mathrm{vol}(E(K_{2}))=0$.

### v2

@@ 8. The bridge number and preorder $\succeq$
The answer to Question 8.1 is positive when $L$ is a Montesinos knot by Corollary 1.7. It is also true when $L^{\prime}$ has at most three bridges, and hence when $G^{\mathrm{orb}}(L^{\prime})$ is finite (by the orbifold theorem).
@@
Question 8.1 is also positive for torus knots.
@@
Let us assume that $K_{1}$ is a $(p_{1},q_{1})$-torus knot and $K_{2}$ a $(p_{2},q_{2})$-torus knot with $3\leq p_{1}<q_{1}$, $3\leq p_{2}<q_{2}$, $p_{1},q_{1}$ coprime, and $p_{2},q_{2}$ coprime. Furthermore, if $p_{i}=3$, then $q_{i}\geq 7$. Since $K_{1}\succeq K_{2}$ there is an epimorphism $\varphi\colon G^{\mathrm{orb}}(K_{1})\twoheadrightarrow G^{\mathrm{orb}}(K_{2})$ which, by the proof of Theorem 1.5(4), sends the infinite cyclic center $Z_{1}$ of $G^{\mathrm{orb}}(K_{1})$ into the infinite cyclic center $Z_{2}$ of $G^{\mathrm{orb}}(K_{2})$. The quotients $G(K_{1})/Z_{1}\cong\pi_{1}^{\mathrm{orb}}(S^{2}(2,p_{1},q_{1}))=T(2,p_{1},q_{1})$ and $G(K_{2})/Z_{2}\cong\pi_{1}^{\mathrm{orb}}(S^{2}(2,p_{2},q_{2}))=T(2,p_{2},q_{2})$ are hyperbolic triangle groups since $\frac{1}{2}+\frac{1}{p_{i}}+\frac{1}{q_{i}}<1$ for $i=1,2$. In particular, they are centerless, and thus the epimorphism $\varphi$ induces an epimorphism $\bar{\varphi}\colon T(2,p_{1},q_{1})\twoheadrightarrow T(2,p_{2},q_{2})$. Like in the proof of Proposition 6.6, it follows from the presentation of the triangle group $T(2,p_{1},q_{1})$ that each image $\bar{\varphi}(x)$, $\bar{\varphi}(y)$ and $\bar{\varphi}(z)$ is not trivial, otherwise the image $\bar{\varphi}(T(2,p_{1},q_{1}))$ would be trivial. Therefore, the non-trivial elements $a=\bar{\varphi}(x)$ and $b=\bar{\varphi}(y)$ verifies $a^{2}=1$ and $b^{p^{\prime}_{1}}=1$ with $p^{\prime}_{1}$ dividing $p_{1}$. The elements $a$ and $b$ generate the image $\bar{\phi}(T(2,p_{1},q_{1}))=T(2,p_{2},q_{2})$ which is a discrete subgroup of $\mathit{PSL}(2,\mathbb{R})$ since $T(2,p_{2},q_{2})$ is a hyperbolic triangle group. As in the proof of Proposition 6.6, up to taking suitable powers $u=a^{\pm 1}$ and $v=b^{k}$, $k$ coprime to $p^{\prime}_{1}$ to normalize the matrix representatives of $u$ and $v$ in $SL(2,\mathbb{R})$, it follows that at least one of Cases (I)–(VII) in [41, Theorem 2.3] holds. Since the triangle group $T(2,p_{2},q_{2})$ is co-compact, Case (II) is impossible by [41, Figures 2 and 3]. Case (III) cannot happens, otherwise $0=\operatorname{tr}(u)=\operatorname{tr}(v)$ and $\bar{\varphi}(T(2,p_{1},q_{1}))=T(2,\ell,n)$ with $\operatorname{tr}(u)=\operatorname{tr}(v)=\cos(\pi-\frac{\pi}{\ell})$ by [41, Section 3]. Then $\ell=2$, which contradicts $\frac{1}{2}+\frac{1}{\ell}+\frac{1}{n}<1$ because $T(2,\ell,n)=T(2,p_{2},q_{2})$ is a hyperbolic triangle group. If Case (IV) happens, then $u$ and $v$ generate a hyperbolic triangle group $T(2,3,n)$ by [41, Section 3]. Since $T(2,3,n)=T(2,p_{2},q_{2})$ with $3\leq p_{2}<q_{2}$ and $p_{2}$ coprime to $q_{2}$, it follows that $p_{2}=3$ and $q_{2}=n$. Therefore,
@@
For a prime link with infinite orbifold group, a positive answer to Question 8.1 would follow from a positive answer to the following question which is a stronger version of the well-known meridional rank conjecture by Cappell and Shaneson [1, Problem 1.11] about whether there is an equality between the bridge number $b(L)$ and the meridional rank $m(L)$ of a link $L$.
Question 8.4.
@@
For a link $L$ as in Question 8.4, $\Sigma_{2}(L)$ is aspherical by Lemma 3.1(3). Therefore, a torsion element in $G^{\mathrm{orb}}(L)$ has order $2$ and is the image of a meridian of $L$.
@@
It follows from results in [2] and [3] that the answer to Question 8.4 is positive for some special classes of links including Montesinos links, twisted links (see [2] for the definition), and arborescent links associated to bipartite trees with even weights. In these articles, it is shown that for such a link, a lower bound on the meridional rank obtained via Coxeter quotients of the link group matches the bridge number of the link. A Coxeter quotient of a link group $G(L)$ is a Coxeter group $C$ with an epimorphism from $G(L)$ onto $C$ which sends the meridians of $G(L)$ to reflections in $C$. The reflection rank of a Coxeter group $C$ is the minimal number of reflections needed to generate $C$. The Coxeter rank $\mathrm{cox}(L)$ of a link $L$ is the maximum of the reflection ranks of the Coxeter quotients of $G(L)$. For a prime link with infinite orbifold group, the inequalities